OpenStudy (anonymous):

Evaluate 2a^2 - 10 ÷ b for a = 3 and b = 5.

OpenStudy (anonymous):

I got the answer 16, can you tell me if I'm right

OpenStudy (anonymous):

its a little bit unclear. is it (2a^2-10)/b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so the answer is 1.6

OpenStudy (anonymous):

How did you get that

OpenStudy (anonymous):

you need to do PEMDS, dont forget that.

OpenStudy (texaschic101):

(2a^2 - 10) / b a = 3 and b = 5 now we will sub (2(3^2) - 10) / 5 (2(9) - 10) / 5 (18 - 10) / 5 8/5 = 1 3/5 or 1.6

OpenStudy (anonymous):

none of the equation is in parentheses. I think you have to dived 10/5 first

OpenStudy (texaschic101):

oh...I thought you said there was parenthesis...let me try this again.. 2a^2 - 10/5 2(3^2) - 10/5 2(9) - 2 18 - 2 = 16

OpenStudy (texaschic101):

If there is no parenthesis, then you are correct